the length of a rectangle is 7 inches longer than twice the width. What are the diemsions of the rec. if its perimeter is 74
Answer by nyc_function(2741)

Let length = 2x + 7


Let width = x


Let perimeter = 74


P = 2L + 2W


74 = 2(2x + 7) + 2x


74 = 4x + 14 + 2x


74 = 6x + 14


74 - 14 = 6x


60 = 6x


60/6 = x


10 = x

The width is 10 inches.

The length is 27 inches.